# No Minima, No Collisions: Combining Modulation and Control Barrier Function Strategies for Feasible Dynamical Collision Avoidance

### v1

@@ II-A Safety-Critical Control for Dynamic Obstacle Avoidance
Both subjects of the work, Mod-DS and CBF-QP, define the notion of robot safety and generate safe control algorithms based on boundary functions. Given a continuously differential function $h$, state $x\in\mathbb{R}^{d}$ and time $t\in\mathbb{R}$, $h:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a boundary function if the safe set $C$ (outside the obstacle), the boundary set $\partial C$ (on the boundary of the obstacle), and the unsafe set $\neg C$ (inside the obstacles) of the system can be defined as in (1), (2), (3[2]. The boundary function $h$ is time-dependent for scenarios where the unsafe set are moving. In static obstacle avoidance, boundary function $h(x,t)=h(x)$.
$C=\{x\in\mathbb{R}^{d}:h(x,t)>0\}$ (1)
$\partial C=\{x\in\mathbb{R}^{d}:h(x,t)=0\}$ (2)
$\neg C=\{x\in\mathbb{R}^{d}:h(x,t)<0\}$ (3)
In practice, $h(x,t)$ is often measured as the distance from the controlled agent to the obstacle surface boundary, i.e. the signed distance function. Given environments defined by boundary functions, the goal of a safety-critical controller is to generate an admissible input $u$ that will ensure the state of the robot $x$ is always within the safe set $C$ and eventually reaches a target state $x^{*}\subset C\in\mathbb{R}^{d}$.
@@ II-C Assumptions
Safe Controller: The safe-controller $u$ takes as input the current state $x$, the nominal control $u_{\text{nom}}$, and outputs the desired safe control input to the system, based on the robot’s knowledge of the environment given by the obstacle boundary function $h(x,t)$. In Section IV and VI, $u_{\text{nom}}$ is defined assuming single integrator agent dynamics, while in Section VIII we formulate $u_{\text{nom}}$ to approximate linear dynamical systems assuming higher order control systems; i.e., Dubin’s car.
Obstacle Shapes: In this work, we analyze and propose safe control methods facing all obstacle geometries, including convex, star-shaped, and non-star-shaped concave obstacles. Star-shaped obstacles are defined as concave obstacles that have a reference point inside the obstacle, from which all rays cross the boundary only once [20]. Without loss of generality to applications in higher-dimensional state space, the method compared and proposed are showcased in 2D navigation and manipulation tasks. We define $x_{\text{rel}}\in\mathbb{R}^{2}$ as the robot position measured in the obstacle based frame, varying based on the geometric center, and orientation of the obstacle. Safe controllers’ abilities in navigating around convex obstacles are tested using a circle with boundary function $h_{\text{conv}}$ defined as in (4), where $c_{r}\in\mathbb{R}^{+}$ is the radius of the circle.
@@
Finally, non-star-shaped obstacles are represented using boundary function $h_{\text{nstar}}$ learned using Gaussian Process Distance Field (GPDF) of a C-shape. Isoline maps showing the change in $h$ values with respect to the relative location of the agent can be found in Fig. 1. $h(x,t)$ functions for convex, star and non-star shapes are constructed to align with the safe set definition in Eq. (1), (2), and (3). Here we limit the scope of our discussion to cases where the size of the obstacles is fixed with no shrinkage, expansion, or deformation. All moving (i.e. time $t$ dependent) obstacles exhibit only translation and rotational motion.
@@ III-A Boundary Function Requirement
According to [2] and [11, 13, 14], CBF-QP and Mod-DS method share the same requirement for the boundary function $h(x,t)$: $h(x,t)$ must be a continuously differentiable function of class $\mathcal{C}^{1}$ satisfying safe set definitions in Eq. (1), (2) and (3).
@@ III-B Control Barrier Functions
In CBF-QP formulation, the boundary function $h(x,t)$ is also called the barrier function. To guarantee safety of the controlled agent, CBF-QP utilizes the Nagumo Set Invariance Theorem.
@@
$C\text{ is set invariant}\iff\dot{h}(x,t)\geq 0\;\forall x\in\partial C\$ (6)
@@
$C\text{ is set invariant}\iff\exists u\;\text{s.t.}\;\dot{h}(x,u,t)\geq-\alpha(h(x,t)).$ (7)
@@
Control barrier functions can be generalized to any nonlinear affine systems of the form in (8), where $x\in\mathbb{R}^{d},u\in\mathbb{R}^{p}$, and $f,g$ are Lipschitz continuous. The CBF conditions in (7) can be used to formulate a quadratic programming problem that guarantees safety by enforcing the set invariance of the safety set $C$ defined in (1). For general control affine systems, CBF-QP is defined as in (9).
@@
$L_{f}h(x,t)+L_{g}h(x,t)u+\frac{\partial h(x,t)}{\partial t}\geq-\alpha(h(x,t))$ (9)
@@
$\nabla_{x}h(x,t)^{\top}u+\frac{\partial h(x,t)}{\partial t}\geq-\alpha(h(x,t))$ (11)
Closed-Form Solution: CBF-QP has closed-form solutions in single obstacle environments when robot input limitations are ignored. Solving the convex optimization problem in (9) and (11) using KKT conditions, the explicit CBF-QP solutions are respectively (12) and (11). For simplicity of the solutions, we abuse notations and let $\alpha=\alpha(h(x,t))+\frac{\partial h(x,t)}{\partial t}$, $L_{f}h=L_{f}h(x,t)$, $L_{g}h=L_{g}h(x,t)$ and $\nabla_{x}h=\nabla_{x}h(x,t)$.
@@ III-C Mod-DS Approach
In Mod-DS, obstacle avoidance is achieved by multiplying a full-rank locally active matrix $M(x,t)\in\mathbb{R}^{d\times d}$ to a nominal DS $\dot{x}_{\text{nom}}$ and reshaping the flow by reducing the relative speed of the robot towards the obstacle, while increasing or maintaining the speed in the directions tangent to the obstacle surface represented by $h(x,t)$. Mod-DS is computed for moving obstacles as [13, 14],
$u_{\text{mod}}=\dot{x}_{\text{mod}}=M(x,t)(\dot{x}_{\text{nom}}-\dot{x}_{o})+\dot{x}_{o}$ (14)
$\dot{x}_{o}=\dot{x}^{o}_{L}+\dot{x}^{o}_{R}\times(x-x_{o}),$ (15)
where $x_{o}$ is an obstacle’s axis of rotation where its motion can be characterized about, $\dot{x}^{o}_{L},\dot{x}^{o}_{R}\in\mathbb{R}^{d}$ are the linear and angular velocity vector of the obstacle with respect to point $x_{o}$. In a static scenario, Eq. (16) reduces to $\dot{x}=M(x)\dot{x}_{\text{nom}}$.
@@
$M(x,t)$ is constructed analogous to an Eigendecomposition, where $E(x,t)\in\mathbb{R}^{d\times d}$ is the basis matrix and $D(x,t)\in\mathbb{R}^{d\times d}$ is the diagonal scaling matrix. The basis matrix $E(x,t)$ consists of a direction vector facing the obstacle $d(x,t)\in\mathbb{R}^{d\times 1}$ and a hyperplane $H(x,t)\in\mathbb{R}^{d\times(d-1)}$ tangent to $h(x,t)$. The hyperplane $H(x,t)$ is an orthonormal basis formed by basis vectors $e_{1}(x,t),...,e_{d-1}(x,t)\in\mathbb{R}^{d\times 1}$. The basis matrix $E(x,t)$ enforces the redistribution of velocities along $d(x,t),e_{1}(x,t),...,e_{d-1}(x,t)$. The magnitudes of the velocities in the redistributed directions are then modified by $D(x,t)$ as in Eq. (16). To achieve multiple obstacle avoidance, a weighted sum of the stretched velocity computed with respect to each obstacle is used, the details of which can be found in [11].
$M(x,t)=E(x,t)D(x,t)E(x,t)^{-1}\\ D(x,t)=\text{diag}(\lambda(x,t),\lambda_{e}(x,t),\dots,\lambda_{e}(x,t))\\ H(x,t)=[e_{1}(x,t)\;...\;e_{d-1}(x,t)]\\ E(x,t)=[d(x,t)\;H(x,t)]\\ \end{aligned}$ (16)
@@
There exist 3 main variants of Mod-DS approaches, different in terms of how $E(x,t)$ and $D(x,t)$ are constructed: normal Mod-DS, reference Mod-DS and on-manifold Mod-DS [11][13][16]. The properties of each method are detailed in Table I. Since any smooth vector field in a sphere world has at least as many topologically critical points as obstacles [21], we consider a method to be able to handle certain types of obstacles if the modified DS has at most 1 local minimum per obstacle given a fully-actuated nominal DS. On any non-equilibrium points in the modified DS, Mod-DS approaches have been proven to ensure impenetrability at the boundary in the sense of von Neuman and if the nominal DS $\dot{x}_{\text{nom}}$ is globally asymptotically stable (g.a.s) then Mod-DS is g.a.s. to the same equilibrium point $x^{*}\in\mathbb{R}^{d}$ [11, 16].
@@
Normal Mod-DS: In normal Mod-DS, vector $n(x,t)$ normal to the boundary function $h(x,t)$ is used as the direction towards the obstacle when constructing the basis matrix $E(x,t)$ (17). It is noteworthy that $\lambda^{*}(x)$ and $\lambda^{*}_{e}(x)$ values listed in Table I are only popular choices of $D(x,t)$ but any $\lambda(x,t)$, $\lambda_{e}(x,t)$ satisfying conditions in (18) will be feasible options. As the earliest proposed Mod-DS variants, normal Mod-DS only guarantees convergence to target $x^{*}$ in convex obstacle avoidance.
$d_{\text{s}}(x,t)=n(x,t)=\frac{\nabla_{x}h(x,t)}{||\nabla_{x}h(x,t)||_{2}}$ (17)
$\lim_{h(x,t)\rightarrow 0}\lambda(x,t)=0\text{ and }\lambda_{e}(x,t)>0\;\;\forall x\in\mathbb{R}^{d}$ (18)
@@
Reference Mod-DS: The only difference between reference Mod-DS and normal Mod-DS is that reference direction $r(x,t)$ is used instead of $n(x,t)$ to form the basis matrix $E(x,t)$ (19). $r(x,t)$ points from the agent’s current location $x$ to a reference point $r^{*}(t)\in\mathbb{R}^{d}$. $r^{*}$ is a constant vector when the unsafe set is static. This change enables the reference Mod-DS to handle star-shaped obstacles, in addition to convex ones.
@@
$r^{*}(t)\in\mathbb{R}^{d}$ is a reference point if and only if it is strictly inside unsafe set $\neg C$, where all rays from which have one and one only intersection with the boundary set $\partial C$. An obstacle or an unsafe set is considered star-shaped if it has a minimum of 1 reference point.
$d_{\text{r}}(x,t)=r(x,t)=\frac{x-r^{*}(t)}{||x-r^{*}(t)||_{2}}$ (19)
Type $D(x)$ $d(x,t)$ Obstacle Local Minima #
Normal $\lambda^{*}(x,t)=1-\frac{1}{h(x,t)+1}$, $\lambda_{e}^{*}(x,t)=1+\frac{1}{h(x,t)+1}$ $n(x,t)$ convex 1
Reference $\lambda^{*}(x,t)=1-\frac{1}{h(x,t)+1}$, $\lambda_{e}^{*}(x,t)=1+\frac{1}{h(x,t)+1}$ $r(x,t)$ convex, star-shaped 1
On-Manifold $D_{\nabla}(x,t)+D_{\phi}(x,t)+\varphi_{h}(x,t)D_{h}(x,t)$ $n(x,t)$ all 0
TABLE I: Properties of 3 main types of Mod-DS, including the definition of their basis matrix and diagonal scaling matrix, and the types of obstacles they can navigate without inducing a trapping region in the modified DS, are documented. For obstacle types that a Mod-DS variant can navigate without being trapped, the number of local minima that the approach induced per obstacle is listed. The definition of $n(x,t)$ and $r(x,t)$ can be found in Eq. (17) and (19)
@@
$d_{\text{onM}}(x,t)=n(x,t)=\frac{\nabla_{x}h(x,t)}{||\nabla_{x}h(x,t)||_{2}}$ (20)
@@
$\lambda(x,t)=\lambda_{\nabla}(x,t)+\lambda_{\phi}(x,t)+\varphi_{h}(x,t)\lambda_{h}(x,t)$ (21)
The technical details of on-manifold Mod-DS can be found in [16]. In this work, we are more interested in the obstacle exit strategy $\phi(x,t)$ used in $\lambda_{\phi}$ to help efficiently circumnavigate both convex and concave unsafe regions. Let $e^{0}$ be one out of $m$ candidate directions in $\mathbb{R}^{d-1}$, subject to $e^{0}\notin\mathcal{N}(H(x,t))$, and $m\geq 2^{d-1}$. A geodesic approximation method can then utilize a first-order approximation to the obstacle surface to closely approximate a path $X=[x^{0},x^{1},...,x^{N}]$ exiting the obstacle on its isosurface, where horizon $N\in\mathbb{N}$ is a natural number [16]. Function $\phi(x,t)\in\mathbb{R}^{d}$ then outputs, among $m$ candidate directions, the one with the smallest associated potential $P_{N}$, where $\beta$ is the step size and $p(x)$ is a user-defined reward function. In most applications, $p(x)$ can be the distance from $x_{i}$ to the target $x^{*}$.
$x^{i+1}$ $=\beta H(x^{i},t)H(x^{i},t)^{\top}e^{i}+x^{i}$ (22)
$e^{i+1}$ $=\frac{H(x^{i},t)H(x^{i},t)^{\top}e^{i}}{||H(x^{i},t)H(x^{i},t)^{\top}e^{i}||_{2}}$

### v2

@@ II-A Safety-Critical Control for Dynamic Obstacle Avoidance
Both subjects of the work, Mod-DS and CBF-QP, define the notion of robot safety and generate safe control algorithms based on boundary functions. Given a continuously differential function $h_{o}$ for an obstacle $o$ in the detected obstacle set $O$, state $x\in\mathbb{R}^{d}$ and obstacle state $x_{o}\in\mathbb{R}^{d^{\prime}}$, $h_{o}:\mathbb{R}^{d}\times\mathbb{R}^{d^{\prime}}\rightarrow\mathbb{R}$ is a boundary function if the safe set $C_{o}$ (outside the obstacle), the boundary set $\partial C_{o}$ (on the boundary of the obstacle), and the unsafe set $\neg C_{o}$ (inside the obstacles) of the system can be defined as in (1), (2), (3[2].
$C_{o}=\{x\in\mathbb{R}^{d},x_{o}\in\mathbb{R}^{d^{\prime}}:h_{o}(x,x_{o})>0\}$ (1)
$\partial C_{o}=\{x\in\mathbb{R}^{d},x_{o}\in\mathbb{R}^{d^{\prime}}:h(x,x_{o})=0\}$ (2)
$\neg C_{o}=\{x\in\mathbb{R}^{d},x_{o}\in\mathbb{R}^{d^{\prime}}:h(x,x_{o})<0\}$ (3)
Note that The boundary function $h_{o}(x,x_{o})$ can be simplified to $h_{o}(x)$ when the obstacle state is static, i.e. $x_{o}$ is constant, and to $h(x,x_{o})$ when the environment contains only one obstacle. In practice, $h_{o}(x,x_{o})$ is often measured as the distance from the controlled agent to the obstacle surface boundary, i.e. the signed distance function. Given environments defined by boundary functions, the goal of a safety-critical controller is to generate an admissible input $u$ that will ensure the state of the robot $x$ is always within the safe set $C$ defined in (4) and eventually reaches a target state $x^{*}\subset C\in\mathbb{R}^{d}$.
$C=\{x\in\mathbb{R}^{d},x_{o}\in\mathbb{R}^{d^{\prime}}:h_{o}(x,x_{o})>0,\forall o\in O\}$ (4)
$\partial C=\{x\in\mathbb{R}^{d},x_{o}\in\mathbb{R}^{d^{\prime}}:h(x,x_{o})=0,\forall o\in O\}$ (5)
$\neg C=\{x\in\mathbb{R}^{d},x_{o}\in\mathbb{R}^{d^{\prime}}:h(x,x_{o})<0,\forall o\in O\}$ (6)
@@ II-C Assumptions
Safe Controller: The safe-controller $u$ takes as input the current state $x$, the nominal control $u_{\text{nom}}$, and outputs the desired safe control input to the system, based on the robot’s knowledge of the environment given by the obstacle boundary functions $h_{o}(x,t),\forall o\in O$. In Section IV and VI, $u_{\text{nom}}$ is defined assuming single integrator agent dynamics, while in Section VIII we formulate $u_{\text{nom}}$ to approximate linear dynamical systems assuming higher order control systems; i.e., Dubin’s car.
Obstacle Shapes: In this work, we analyze and propose safe control methods facing all obstacle geometries, including convex, star-shaped, and non-star-shaped concave obstacles. Star-shaped obstacles are defined as obstacles that have a reference point inside the obstacle, from which all rays cross the boundary only once [20]. Without loss of generality to applications in higher-dimensional state space, the methods compared and proposed are showcased in 2D navigation and manipulation tasks. We define $x_{\text{rel}}(x,x_{o})\in\mathbb{R}^{2}$ as the robot position measured in the obstacle-based frame, which changes with the obstacle’s geometric center and orientation described in $x_{o}$. Safe controllers’ abilities in navigating around convex obstacles are tested using a circle with boundary function $h_{\text{conv}}$ defined as in (7), where $c_{r}\in\mathbb{R}^{+}$ is the radius of the circle.
@@
Finally, non-star-shaped obstacles are represented using boundary function $h_{\text{nstar}}$ learned using the Gaussian Process Distance Field (GPDF) of a C-shape. Isoline maps showing the change in $h_{\text{nstar}}$ values with respect to the relative location of the agent can be found in Fig. 1. Boundary functions for convex, star, and non-star shapes are constructed to align with the safe set definition in Eq. (1), (2), and (3). Here we limit the scope of our discussion to cases where the size of the obstacles is fixed with no shrinkage, expansion, or deformation. All moving obstacles exhibit only translation and rotational motion, i.e.
$x_{o}=\begin{bmatrix}x_{L}^{o}\\ x_{R}^{o}\\ \omega_{o}\end{bmatrix}$ (9)
$R(x_{o})=\exp{(\hat{\omega}_{o}x_{R}^{o})}$ (10)
$x_{\text{rel}}(x,x_{o})=R(x_{o})^{\top}(x-x_{L}^{o})$ (11)
where $x_{L}^{o}$ is the obstacle’s point of rotation on the rotation axis $\omega_{o}$, $x_{R}^{o}$ is the rotation angle, and $R=R(x_{R}^{o})$ is the transformation matrix from the obstacle frame to the world frame, given axis-angle representation of $x_{L}^{o}$ and $x_{R}^{o}$. $\dot{x}_{L}^{o}$ is linear velocity, and $\dot{x}_{R}^{o}$ is the angular velocity of the obstacle. Substituting Eq. (11) into boundary functions, it can be deduced that
$h_{o}(x,x_{o})=h_{o}(x_{\text{rel}})=h_{o}(R(x_{R}^{o})^{\top}(x-x_{L}^{o}))$ (12)
@@ III-A Boundary Function Requirement
According to [2] and [11, 13, 14], CBF-QP and Mod-DS method share the same requirement for the boundary function $h_{o}(x,x_{o})$: $h_{o}(x,x_{o})$ must be a continuously differentiable function of class $\mathcal{C}^{1}$ satisfying safe set definitions in Eq. (1), (2) and (3). In the background section, we review CBF-QP and Mod-DS method in single obstacle avoidance settings, where $h_{o}(x,x_{o})=h(x,x_{o})$.
@@ III-B Control Barrier Functions
In CBF-QP formulation, the boundary function $h(x,x_{o})$ is also called the barrier function. To guarantee safety of the controlled agent, CBF-QP utilizes the Nagumo Set Invariance Theorem.
@@
$C\text{ is set invariant}\iff\dot{h}(x,x_{o})\geq 0\;\forall x\in\partial C\$ (13)
@@
$C\text{ is set invariant}\iff\exists u\;\text{s.t.}\;\dot{h}(x,x_{o},\dot{x}_{o},u)\geq-\alpha(h(x,x_{o})).$ (14)
@@
Control barrier functions can be generalized to any nonlinear affine systems of the form in (15), where $x\in\mathbb{R}^{d},x_{o}\in\mathbb{R}^{d^{\prime}},u\in\mathbb{R}^{p}$, and $f,g$ are Lipschitz continuous. The CBF condition in (14) can be used to formulate a quadratic programming problem that guarantees safety by enforcing the set invariance of the safety set $C_{o}$ defined in (1). For general control affine systems, CBF-QP is defined as in (16).
@@
$L_{f}h(x,x_{o})+L_{g}h(x,x_{o})u+\nabla_{x_{o}}h(x,x_{o})\dot{x}_{o}\geq-\alpha(h(x,x_{o}))$ (16)
@@
$\nabla_{x}h(x,x_{o})^{\top}u+\nabla_{x_{o}}h(x,x_{o})\dot{x}_{o}\geq-\alpha(h(x,x_{o}))$ (18)
Closed-Form Solution: CBF-QP has closed-form solutions in single obstacle environments when robot input limitations are ignored. Solving the convex optimization problem in (16) and (18) using KKT conditions, the explicit CBF-QP solutions are respectively (19) and (18). For simplicity of the solutions, we abuse notations and let $\alpha=\alpha(h(x,x_{o}))+\nabla_{x_{o}}h(x,x_{o})\dot{x}_{o}$, $L_{f}h=L_{f}h(x,x_{o})$, $L_{g}h=L_{g}h(x,x_{o})$ and $\nabla_{x}h=\nabla_{x}h(x,x_{o})$.
@@ III-C Mod-DS Approach
In Mod-DS, obstacle avoidance is achieved by multiplying a full-rank locally active matrix $M(x,x_{o})\in\mathbb{R}^{d\times d}$ to a nominal DS $\dot{x}_{\text{nom}}$ and reshaping the flow by reducing the relative speed of the robot towards the obstacle, while increasing or maintaining the speed in the directions tangent to the obstacle surface represented by $h(x,x_{o})$. Mod-DS is computed for moving obstacles as [13, 14],
$u_{\text{mod}}=\dot{x}_{\text{mod}}=M(x,x_{o})(\dot{x}_{\text{nom}}-\bar{\dot{x}}_{o})+\bar{\dot{x}}_{o}$ (21)
$\bar{\dot{x}}_{o}=\dot{x}^{o}_{L}+\dot{x}^{o}_{R}\times(x-x_{L}^{o}),$ (22)
where $x^{o}_{L}$, $\dot{x}^{o}_{L}$, and $\dot{x}^{o}_{R}\in\mathbb{R}^{d}$ are the defined as in Eq. (9). In a static scenario, Eq. (23) reduces to $\dot{x}=M(x)\dot{x}_{\text{nom}}$.
@@
$M(x,x_{o})$ is constructed analogous to an Eigendecomposition, where $E(x,x_{o})\in\mathbb{R}^{d\times d}$ is the basis matrix and $D(x,x_{o})\in\mathbb{R}^{d\times d}$ is the diagonal scaling matrix. The basis matrix $E(x,x_{o})$ consists of a direction vector facing the obstacle $d(x,x_{o})\in\mathbb{R}^{d\times 1}$ and a hyperplane $H(x,x_{o})\in\mathbb{R}^{d\times(d-1)}$ tangent to $h(x,x_{o})$. The hyperplane $H(x,x_{o})$ is an orthonormal basis formed by basis vectors $e_{1}(x,x_{o}),...,e_{d-1}(x,x_{o})\in\mathbb{R}^{d\times 1}$. The basis matrix $E(x,x_{o})$ enforces the redistribution of velocities along $d(x,x_{o}),e_{1}(x,x_{o}),...,e_{d-1}(x,x_{o})$. The magnitudes of the velocities in the redistributed directions are then modified by $D(x,x_{o})$ as in Eq. (23). To achieve multiple obstacle avoidance, a weighted sum of the stretched velocity computed with respect to each obstacle is used, the details of which can be found in [11].
$M(x,x_{o})=E(x,x_{o})D(x,x_{o})E(x,x_{o})^{-1}\\ D(x,x_{o})=\text{diag}(\lambda(x,x_{o}),\lambda_{e}(x,x_{o}),\dots,\lambda_{e}(x,x_{o}))\\ H(x,x_{o})=[e_{1}(x,x_{o})\;...\;e_{d-1}(x,x_{o})]\\ E(x,x_{o})=[d(x,x_{o})\;H(x,x_{o})]\\ \end{aligned}$ (23)
@@
There exist 3 main variants of Mod-DS approaches, different in terms of how $E(x,x_{o})$ and $D(x,x_{o})$ are constructed: normal Mod-DS, reference Mod-DS and on-manifold Mod-DS [11][13][16]. The properties of each method are detailed in Table I. Since any smooth vector field in a sphere world has at least as many topologically critical points as obstacles [21], we consider a method to be able to handle certain types of obstacles if the modified DS has at most 1 local minimum per obstacle given a fully-actuated nominal DS. Mod-DS approaches have been proven to ensure impenetrability at the obstacle boundary in the sense of von Neuman and if the nominal DS $\dot{x}_{\text{nom}}$ is globally asymptotically stable (g.a.s) then the modulated DS is g.a.s. to the same target point $x^{*}\in\mathbb{R}^{d}$ on all non-equilibrium points [11, 16].
@@
Normal Mod-DS: In normal Mod-DS, vector $n(x,x_{o})$ normal to the boundary function $h(x,x_{o})$ is used as the direction towards the obstacle when constructing the basis matrix $E(x,x_{o})$ (24). It is noteworthy that $\lambda^{*}(x,x_{o})$ and $\lambda^{*}_{e}(x,x_{o})$ values listed in Table I are only popular choices of $D(x,x_{o})$ but any $\lambda(x,x_{o})$, $\lambda_{e}(x,x_{o})$ satisfying conditions in (25) will be feasible options. As the earliest proposed Mod-DS variants, normal Mod-DS only guarantees convergence to target $x^{*}$ in convex obstacle avoidance.
$d_{\text{n}}(x,x_{o})=n(x,x_{o})=\frac{\nabla_{x}h(x,x_{o})}{||\nabla_{x}h(x,x_{o})||_{2}}$ (24)
$\lim_{h(x,x_{o})\rightarrow 0}\lambda(x,x_{o})=0\text{ and }\lambda_{e}(x,x_{o})>0\;\;\forall x\in\mathbb{R}^{d}$ (25)
@@
Reference Mod-DS: The only difference between reference Mod-DS and normal Mod-DS is that reference direction $r(x,t)$ is used instead of $n(x,x_{o})$ to form the basis matrix $E(x,x_{o})$ (26). $r(x,x_{o})$ points from the agent’s current location $x$ to a reference point $r^{*}(x_{o})\in\mathbb{R}^{d}$. $r^{*}$ is a constant vector when the unsafe set is static. This change enables the reference Mod-DS to handle star-shaped obstacles, in addition to convex ones.
@@
$r^{*}(x_{o})\in\mathbb{R}^{d}$ is a reference point if and only if it is strictly inside unsafe set $\neg C_{o}$, where all rays from which have one and one only intersection with the boundary set $\partial C_{o}$. An obstacle or an unsafe set is considered star-shaped if it has a minimum of one reference point.
$d_{\text{r}}(x,x_{o})=r(x,x_{o})=\frac{x-r^{*}(x_{o})}{||x-r^{*}(x_{o})||_{2}}$ (26)
Type $D(x,x_{o})$ $d(x,x_{o})$ Obstacle Local Minima #
Normal $\lambda^{*}(x,x_{o})=1-\frac{1}{h(x,x_{o})+1}$, $\lambda_{e}^{*}(x,x_{o})=1+\frac{1}{h(x,x_{o})+1}$ $n(x,x_{o})$ convex 1
Reference $\lambda^{*}(x,x_{o})=1-\frac{1}{h(x,x_{o})+1}$, $\lambda_{e}^{*}(x,x_{o})=1+\frac{1}{h(x,x_{o})+1}$ $r(x,x_{o})$ convex, star-shaped 1
On-Manifold $D_{\nabla}(x,x_{o})+D_{\phi}(x,x_{o})+\varphi_{h}(x,x_{o})D_{h}(x,x_{o})$ $n(x,x_{o})$ all 0
TABLE I: Properties of 3 main types of Mod-DS, including the definition of their basis matrix and diagonal scaling matrix, and the types of obstacles they can navigate without inducing a trapping region in the modified DS, are documented. For obstacle types that a Mod-DS variant can navigate without being trapped, the number of local minima that the approach induced per obstacle is listed. The definition of $n(x,x_{o})$ and $r(x,x_{o})$ can be found in Eq. (24) and (26)
@@
$d_{\text{onM}}(x,x_{o})=n(x,x_{o})=\frac{\nabla_{x}h(x,x_{o})}{||\nabla_{x}h(x,x_{o})||_{2}}$ (27)
@@
$\lambda(x,x_{o})=\lambda_{\nabla}(x,x_{o})+\lambda_{\phi}(x,x_{o})+\varphi_{h}(x,x_{o})\lambda_{h}(x,x_{o})$ (28)
The technical details of on-manifold Mod-DS can be found in [16]. In this work, we are more interested in the obstacle exit strategy $\phi(x,x_{o})$ used in $\lambda_{\phi}$ to help efficiently circumnavigate both convex and concave unsafe regions. Let $e^{0}$ be one out of $m$ candidate directions in $\mathbb{R}^{d-1}$, subject to $e^{0}\notin\mathcal{N}(H(x,x_{o}))$, and $m\geq 2^{d-1}$. A geodesic approximation method can then utilize a first-order approximation to the obstacle surface to closely approximate a path $X=[x^{0},x^{1},...,x^{N}]$ exiting the obstacle on its isosurface, where horizon $N\in\mathbb{N}$ is a natural number [16]. Function $\phi(x,x_{o})\in\mathbb{R}^{d}$ then outputs, among $m$ candidate directions, the one with the smallest associated potential $P_{N}$, where $\beta$ is the step size and $p(x)$ is a user-defined reward function. In most applications, $p(x)$ can be the distance from $x_{i}$ to the target $x^{*}$.
$x^{i+1}$ $=\beta H(x^{i},x_{o})H(x^{i},x_{o})^{\top}e^{i}+x^{i}$ (29)
$e^{i+1}$ $=\frac{H(x^{i},x_{o})H(x^{i},x_{o})^{\top}e^{i}}{||H(x^{i},x_{o})H(x^{i},x_{o})^{\top}e^{i}||_{2}}$